OpenStudy (anonymous):

Can you help me with this problem? Weights of male mountain lions follow the normal distribution with a median of 150 lb and an interquartile range of 8.2 lb. 1. Find the 75th percentile of the weights. 2. Find the 95th percentile of the weights.

OpenStudy (anonymous):

The standard deviation can be calculated using the formula σ=Interquartile range/1.34896 For this case the standard deviation is 8.2/1.34896= 6.079 A standard normal distribution table shows the z-score for a cumulative probability of 75% is z = 0.675 The median and the mean are the same. Therefore the following formula can be used: z=X−μσ Substituting we get 0.675=X−1506.079 Solving for X gives X = 150 + 4.1 Can you do the next one?

OpenStudy (anonymous):

Apply my formula and get the answer to the 2nd question, I cannot give you the full answer else I may be reported.

OpenStudy (anonymous):

It's the formula.

OpenStudy (anonymous):

Mhm.

OpenStudy (kropot72):

A standard normal distribution table shows the z-score for a cumulative probability of 95% is z = 1.645 The median and the mean are the same. Therefore the following formula can be used: \[z=\frac{X-\mu}{\sigma}\] Substituting we get \[1.645=\frac{X-150}{6.079}\] \[X=150+(1.645\times 6.079)=you\ can\ finish\]

OpenStudy (kropot72):

@Dharmaputra_993 Note the mistakes in your reproduction of the solution to 1.

OpenStudy (anonymous):

Didn't see those before, sorry and thanks!